OpenStudy (anonymous):

An alloy of tin is 14% tin and weighs 22 pounds. A second alloy is 9% tin. How much of the second alloy must be added to the first alloy to get a 12% mixture.

OpenStudy (paki):

a1: 14% tin, 22 pounds a2: 9% tin, X pounds mix: 12% tin, (22 + X) pounds Set the portions of tin before mixing equal to the portion of tin after mixing: (14)(22) + (9)(X) = (12)(22 + X) Solve the equation for X. 308 + 9X = 264 + 12X 44 + 9X = 12X 44 = 3X X = 44/3 pounds = 14 and 2/3 pounds

OpenStudy (yanasidlinskiy):

Set the proportions.